OpenStudy (anonymous):

Lena has been offered a job with two salary options. Option 1 starts at $31,000 and has a $500 raise each year. Option 2 starts at $29,000 and has a 4% raise each year. Which pays more after 5 years? Answer Option 1, at $33,550 Option 2, at $33,550 Option 1, at $35,282 Option 2, at $35,282

OpenStudy (rsadhvika):

Option 1, her salary increases 500 each year. So its arithmetic sequence with a = 31000, d = 500 can you find 5th term ?

OpenStudy (anonymous):

the 5th term is 33000

OpenStudy (rsadhvika):

one small catch here, since you want to know salary after 5 years, you will have to find 6th term. NOT 5th term

OpenStudy (anonymous):

so 33500

OpenStudy (rsadhvika):

Correct. Option 1 : 33,500 -------------------------------(1)

OpenStudy (rsadhvika):

lets find 6th term for Option2 also

OpenStudy (anonymous):

but thats not a choice

OpenStudy (anonymous):

29020

OpenStudy (rsadhvika):

Option 2 starts at $29,000 and has a 4% raise each year. a = 29,000 r = 1.04 \(a_n = a r^{n-1}\) \(a_6 = 29000 (1.04)^{6-1}\) = ?

OpenStudy (anonymous):

29005

OpenStudy (rsadhvika):

check again

OpenStudy (anonymous):

35282?

OpenStudy (rsadhvika):

Yes

OpenStudy (rsadhvika):

Option 2, at $35,282

OpenStudy (anonymous):

thanks for the help

OpenStudy (rsadhvika):

yw